OpenStudy (anonymous):

Solve tan x + 1 = 0 for x.

OpenStudy (anonymous):

start with \[\tan(x)=-1\] and then think of points on the unit circle where sine and cosine are the same number except of opposite sign

OpenStudy (anonymous):

btw i sent a link with the numerical answer to your last question

OpenStudy (anonymous):

tan x + 1 = 0 arc tan (0) = 0 so x+1 = 0 x = -1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

tan a = b arc tan b = a So a = x +1 b = 0 tan x + 1 = 0 arc tan (0) = 0 x+1 = 0 x = -1

OpenStudy (anonymous):

give a general formual for tan x= -1

OpenStudy (kira_yamato):

tan x = -1 x = arctan (-1) = πn-(π/4), n is integer

OpenStudy (anonymous):

x = −π/4 +2nπ